• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Bailey

    Log entry time 22:06:39 on July21,2003

    Entry number 105950

    keyword=target update

    16:30 Turned off transverse lasers
    Rotated holding field from 270 degrees to 0 degrees
    moved target to pick-up coils
    16:35 Waiting for temperatures to stabilize
    16:47 NMR measurement : 40.99
    Turned on longitudinal lasers (4,6,7)
    EPR measurement : 40.04
    17:10 Moved target to empty reference cell
    18:00 pumped reference cell to 140 psig with nitrogen (initial reading -0.14 psig)